OpenStudy (sasogeek):

what's the easiest way to determine whether the binary operation * is associative, if * is defined by \(x*y=\frac{x+y}{1+xy} \)

OpenStudy (anonymous):

Define associativity

OpenStudy (sasogeek):

x*(y*z)=(x*y)*z ... if it's associative, then the right hand side should be equal to the left hand side

OpenStudy (anonymous):

It is associative i did it by verifying LHS=RHS Although intuition pointed it right away

OpenStudy (sasogeek):

well is that the easiest way...? i was thinking maybe there's a faster method

OpenStudy (mr.math):

It clearly follows from the associativity of addition and multiplication.

OpenStudy (sasogeek):

what do u mean by that?

OpenStudy (mr.math):

I mean \((x*y)*z=\frac{(x+y)+z}{1+(xy)z}=\frac{x+(y+z)}{1+x(yz)}=x*(y*z).\)

OpenStudy (sasogeek):

ohhh okay, well i was doing some weird workings but the answer was quite obvious right from the beginning :) thanks

OpenStudy (mr.math):

You're welcome.